# Rethinking the Starting Point: Collaborative Pre-Training for Federated Downstream Tasks

### v3

@@ Appendix A Key Applications
In our experimental comparison, we consider a FL baseline, FedMeta [7], which also incorporates meta-learning. We would like to emphasize that the distinction between our CoPreFL and FedMeta lies in the application, which requires us to conduct meta-update in a totally different way. FedMeta is designed to offer personalization to individual clients, e.g., when a specific client is interested in predicting only diseases A and B, or when a specific self-driving car is interested in the model tailored to a specific weather. In contrast, our emphasis is on creating an initial model that can construct a good “global model” during downstream instead of “personalized models”, targeting the aforementioned applications. This is the reason why we need to update the temporary global model instead of the local models, which is the key technical difference with FedMeta. Another technical difference is the consideration of performance balance in our method. These two key techniques enables CoPreFL to construct a robust initial model that can quickly adapt to “any group of clients” (instead of individual clients) to construct a global model during downstream tasks.
@@ Appendix B Detailed Procedure for CoPreFL in Scenario II
1
21:  Input: $M$ clients in the pre-training phase, with each client $i$ holding their own dataset $D^{p}_{i}$; the server also holds a dataset $D^{s}$.
@@
36:        Downloads $\Phi^{t-1}$ from the server
@@
8:           $\ell^{e}_{D^{p}_{j}}(\Phi^{t})\leftarrow\frac{1}{|D^{p}_{j}|}\sum_{(x,y)\in D^{p}_{j}}\ell(\Phi^{t,e}_{j}(x),y)$// Compute local support loss $\ell^{e}_{D^{p}_{j}}$}
49:           $\Phi^{t,e}_{j}$leftarrow $\Phi^{t,e-1}_{j}$- $\eta\nabla\ell^{e}_{D^{p}_{j}}(\Phi^{t})$// Perform SGD local update using support loss}
510:        end for
@@
612:     $\overline{\Phi^{t}}$leftarrow $\sum_{j\in m}\frac{|D^{p}_{j}|}{\sum_{i\in m}|D^{p}_{i}|}\Phi^{t,E}_{j}$// Model aggregation to construct temporary global model}
713:     for Each split server’s dataset $D^{s}_{j}$ in parallel, Server do
14:        $\mathcal{L}_{D^{s}_{j}}(\overline{\Phi^{t}})$leftarrow $\frac{1}{|D^{s}_{j}|}$ $\sum_{(x,y)\in D^{s}_{j}}{}\ell(\overline{\Phi^{t}}(x),y)$// Server’s loss corresponding to each partition}
@@
817:     Variance across server meta-losses: $\sigma_{D^{s}}^{2}(\overline{\Phi^{t}})=\frac{1}{|m|}\sum_{j\in m}(\mathcal{L}_{D^{s}_{j}}(\overline{\Phi^{t}})-\frac{1}{|m|}\mathcal{L}_{D^{s}}(\overline{\Phi^{t}}))^{2}$
918:     Customized server meta-loss: $\mathcal{L}_{meta}(\overline{\Phi^{t}})=\gamma\mathcal{L}_{D^{s}}(\overline{\Phi^{t}})+(1-\gamma)\sigma_{D^{s}}^{2}(\overline{\Phi^{t}})$
19:     $\Phi^{t}$leftarrow $\overline{\Phi^{t}}$- $\zeta\nabla\mathcal{L}_{meta}(\overline{\Phi^{t}})$// Model meta-updates using customized loss}
1020:  end for
@@ C.1 Dataset Details
In the IID setup, data samples from each class are distributed equally to $|M|=100$ clients for pre-training and $|G|=10$ clients for downstream FL task. Taking the CIFAR-100 dataset in the IID pre-training phase as an example, there are $|M|=100$ clients, each holding 456 images. We will randomly select $|m|$ clients to participate in pre-training in each round. In the IID downstream phase of CIFAR-100, there are $|G|=10$ clients, each holding 300 images. In the non-IID setup, samples within each class are partitioned among $|M|$ and $|G|$ clients using a Dirichlet($\alpha$) distribution for pre-training and downstream task, respectively, with $\alpha=0.5$ selected as is in the literature [33, 28].
In addition to the label distributional shifts discussed in our manuscript, we also explore our method’s capability to handle domain shifts. To assess this, we conducted an additional experiment using the PACS dataset [27]. As PACS dataset consists of four domains (Art, Cartoon, Photo, Sketch), we separate them into three domains for pre-training and one domain for downstream FL, following the one-domain-leave-out setup [30, 47].
@@ C.2 Hyperparameters and Compute Settings
We use ResNet-18 as the model structure for image classification, following the setting of [34, 8]. For our method, the SGD optimizer with a learning rate of $\eta=10^{-3}$ and $\zeta=10^{-3}$ is adopted for both local and meta updates. Both local and meta learning rates are searched within the range of [1e-2, 5e-3, 1e-3, 5e-4]. We searched for learning rates within the range of [1e-2, 5e-3, 1e-3, 5e-4] for local training of all FL pre-training baselines and selected 1e-3 as the optimal learning rate for them. In scenario II, each FL baseline will continue to conduct a few SGD iterations using the server’s data after constructing their global model. We searched for learning rates in the range of [1e-2, 1e-3] for this additional training and selected 1e-3 as the optimal learning rate for the server. Regarding hyperparameters in the q-FFL baseline, we conducted experiments with q-values of 1, 3, and 5 and reported the corresponding best statistics. We select a learning rate $\eta$ from the range [1e-2, 5e-3, 1e-3, 5e-4] for local updates in our CoPreFL and determined that 1e-3 provides the best results. Additionally, for meta-updates in both scenarios, we search for the learning rate $\zeta$ within the range [1e-2, 1e-3] and find that 1e-3 is the optimal value. In the case of the centralized baseline mentioned in Section 4.2, we searched for the optimal learning rate within the range [1e-2, 5e-3, 1e-3, 5e-4, 1e-4], ultimately selecting 1e-3. We utilized the SGD optimizer for all updates across all methods, and the batch size is set to be 32 for all experiments.
@@ D.1 Downstream FL Results with Scenario I Pre-training
Comparison of IID and non-IID downstream tasks. In some cases, when comparing IID and non-IID downstream tasks, we see that data heterogeneity has only a slight impact on the proposed approach. To be more specific, the performance of non-IID downstream FL experiences only a slight performance drop compared to IID downstream FL. In an IID setting, the model must classify all classes present in the system, as each client will have most classes in its local dataset. However, in a non-IID downstream scenario, the model may only need to classify a few classes since each client typically has a limited subset of classes. This simplification of local tasks in the non-IID setting reduces the complexity of evaluation for each client, minimizing the impact of data heterogeneity on the overall federated learning process. This phenomenon aligns with findings from other FL studies, where global model performance is assessed using local test sets from individual clients [29, 12].
@@ D.2 Downstream FL Results with Scenario II Pre-training
Additional results with varying degrees of IIDness and numbers of participants. This section provides supplementary results for scenario II, where the server holds a small portion of the dataset during the pre-training phase. We also consider varying numbers of participants $|m|$, specifically 15, 20, 25, and 30 out of 100 clients, during the pre-training phase for these models. Tables 25, 26, 27, and 28 display the average performance across 10 IID FL downstream tasks and Tables 29, 30, 31, and 32 show the average performance across 10 non-IID FL downstream tasks. In both cases, the downstream FL were initialized by pre-trained models trained on 15, 20, 25, and 30 participants out of 100 clients, respectively, on the CIFAR-100 dataset. For the Tiny-ImageNet dataset, Tables 33, 34, 35, and 36 show the average performance across 10 IID FL downstream tasks and Tables 37, 38, 39, and 40 display the average performance across 10 non-IID FL downstream tasks. In both cases, the downstream FL were also initialized by pre-trained models trained on 15, 20, 25, and 30 participants out of 100 clients, respectively. It is important to note that in this scenario, FedAvg, FedMeta, and q-FFL undergo further training using server data through the SGD optimizer after each method completes its local iterations and obtains its respective global model in each round [43, 3]. Similarly, CoPreFL-SGD is trained using server data with the SGD optimizer on $\Phi^{t}$ in line 17 of Algorithm 1 in each round. This process involves conducting meta-updates and balancing performance and variance using clients’ data first, followed by updating the aggregate model again using the server’s dataset. Finally, CoPreFL follows Algorithm 2, utilizing server data for meta-updates. By incorporating meta-updates using server data to align with our objectives in (3), our pre-training method consistently outperforms other baselines, leading to improved average accuracy and reduced variance. Comparing CoPreFL with CoPreFL-SGD strongly suggests that, rather than conducting a few SGD iterations using server data, which may dilute our objectives, we recommend building pre-training objectives upon server data using meta-updates.
@@
Alternative implementation for baselines in scenario II. In addition to the hybrid training approach introduced in [43, 3], which utilizes clients’ data ($\bigcup_{i\in M}D^{p}_{i}$) to train local models and then refines the aggregated global model using the server’s data ($D^{S}$), we explore an alternative implementation for other FL baselines in scenario II. In this case, we distribute the entire training dataset ($\bigcup_{i\in M}D^{p}_{i}+D^{S}$) to $|M|=100$ clients and select $|m|=20$ participants in each round for federated learning without a further refining step since there is no server’s data in this case. Therefore, each client holds more samples compared to their previous scheme. Table 41 shows the average performance of downstream non-IID FedAvg using non-IID federated methods as initialization. It is important to note that we maintain a fixed data splitting setup for our method, meaning we use $\bigcup_{i\in M}D^{p}_{i}$ for local training and meta-update the temporary model using $D^{S}$. These comparisons also show the superiority of our method.
@@ D.4 Details and Additional Results for FEMNIST Dataset
We also consider the FEMNIST dataset, widely used in FL research, following the data partition provided in [35]. We divide the 62 classes into 52 alphabet classes for the pre-training phase, reserving the remaining 10 digit classes for downstream FL tasks. Instead of using a ResNet-18 model, we employ a model consisting of two 3$\times$3 convolutional layers followed by two linear layers. We fixed the total number of clients as $|M|=100$ for pre-training and $|G|=10$ for downstream FL tasks. During the pre-training phase, we set the number of participants $|m|=20$ and the federated round $T=50$ for each federated pre-trained method. We use the SGD optimizer with a learning rate of $10^{-3}$ and batch size 32 for baselines and our method.
@@ D.5 Implementation Details and Additional Results for Different Initialization Methods
For FL baselines, we additionally consider SCAFFOLD [23], which addresses partial client sampling, FedDyn [1], designed to tackle non-IID issues, and PerFedAvg [15], aiming to provide an adaptable personalized model, for a detailed comparison. We train all FL algorithms for 50 rounds with $|m|=20$ participants selected from $|M|=100$ clients under a non-IID setting (Dirichlet $\alpha=0.5$) , and the final global model is used as initialization for downstream FedAvg. In the case of non-IID related FL, FedDyn, we set the parameter $\alpha$ to 0.01. For personalized FL, PerFedAvg, we employ a two-step gradient descent for local client training introduced in their paper. We use the SGD optimizer with a learning rate of $10^{-3}$, a batch size of 32, and 50 federated rounds for these FL-based baselines.
@@ D.6 Implementation Details and Additional Results for Different Downstream FL Tasks
In addition to the general downstream FL tasks built by FedAvg, we consider FedProx [40] and q-FFL [29], more advanced FL algorithms that addresses heterogeneity and performance balance compared to FedAvg, to examine the robustness and generalizability of our pre-trained method. The experiments are conducted using the CIFAR-100 dataset under non-IID pre-training scenario I. Two additional FL algorithms, non-IID FedProx and non-IID q-FFL, are considered for downstream phase. We randomly sample 5 classes from the 20 available in our CIFAR-100 downstream dataset for each downstream task. The sampled data is then distributed to 10 clients, and the training consists of 50 rounds with 5 local iterations per round, utilizing an SGD optimizer with a learning rate of $10^{-3}$. We set the parameters $\mu=1$ for the proximal term coefficient in FedProx and $q=2$ for the loss-reweighting coefficient in q-FFL, following the optimal values reported by the authors for the CIFAR-100 dataset.
@@ D.7 Implementation Details when A Public Large-Scale Dataset is Available
In addition to utilizing CIFAR-100 and Tiny-ImageNet datasets, where we partition the datasets for pre-training and downstream tasks, we also explore a scenario where public large datasets are available for pre-training phase. We conducted experiments using pre-trained models with the ImageNet dataset [10], a widely used large public dataset for image classification. We sampled 200 images for each of the 1,000 classes in ImageNet_1K as a pre-training dataset. We conducted pre-training using both the centralized method and our proposed CoPreFL with ImageNet_1K. Subsequently, we conducted 10 non-IID FedAvg tasks using the CIFAR-100 dataset and initialized the models with these pre-trained models. For the centralized model in pre-training phase, we trained the model with the SGD optimizer and a learning rate of 1e-3, training the model for 50 epochs. For our proposed method during pre-training, we distributed all the sampled data across $|M|=100$ clients based on non-IID distribution (Dirichlet $\alpha=0.5$), sampling $|m|=20$ clients in each round, and conducted CoPreFL for 50 rounds. Since the goal of this experiment is to demonstrate that even with a centrally-stored public large dataset, we can intentionally distribute the dataset and apply our method, we only consider conducting our method under scenario I. For the downstream FL phase, we apply 10 IID and non-IID FedAvg tasks to the 20-class and 40-class downstream datasets we used in CIFAR-100 and Tiny-ImageNet, respectively. Each task executes FedAvg for $R=10$ rounds using the SGD optimizer with a learning rate of $10^{-3}$. Note that all classes observed during downstream tasks are the seen classes that have already appeared during pre-training, given that CIFAR-100 and Tiny-ImageNet are the subsets of ImageNet.

### v4

@@ Appendix A Key Applications
In our experimental comparison, we consider a FL baseline, FedMeta (Chen et al. 2018), which also incorporates meta-learning. We would like to emphasize that the distinction between our CoPreFL and FedMeta lies in the application, which requires us to conduct meta-update in a totally different way. FedMeta is designed to offer personalization to individual clients, e.g., when a specific client is interested in predicting only diseases A and B, or when a specific self-driving car is interested in the model tailored to a specific weather. In contrast, our emphasis is on creating an initial model that can construct a good “global model” during downstream instead of “personalized models”, targeting the aforementioned applications. This is the reason why we need to update the temporary global model instead of the local models, which is the key technical difference with FedMeta. Another technical difference is the consideration of performance balance in our method. These two key techniques enables CoPreFL to construct a robust initial model that can quickly adapt to “any group of clients” (instead of individual clients) to construct a global model during downstream tasks.
@@ Appendix B Detailed Procedure for CoPreFL in Scenario II
1:  Input: $M$ clients in the pre-training phase, with each client $i$ holding their own dataset $D^{p}_{i}$; the server also holds a dataset $D^{s}$.
@@
6:        Downloads $\Phi^{t-1}$ from the server
@@
8:           $\ell^{e}_{D^{p}_{j}}(\Phi^{t})\leftarrow\frac{1}{|D^{p}_{j}|}\sum_{(x,y)\in D^{p}_{j}}\ell(\Phi^{t,e}_{j}(x),y)$ Compute local support loss $\ell^{e}_{D^{p}_{j}}$}
9:           $\Phi^{t,e}_{j}$leftarrow $\Phi^{t,e-1}_{j}$- $\eta\nabla\ell^{e}_{D^{p}_{j}}(\Phi^{t})$ Perform SGD local update using support loss}
10:        end for
@@
12:     $\overline{\Phi^{t}}$leftarrow $\sum_{j\in m}\frac{|D^{p}_{j}|}{\sum_{i\in m}|D^{p}_{i}|}\Phi^{t,E}_{j}$ Model aggregation to construct temporary global model}
13:     for Each split server’s dataset $D^{s}_{j}$ in parallel, Server do
14:        $\mathcal{L}_{D^{s}_{j}}(\overline{\Phi^{t}})$leftarrow $\frac{1}{|D^{s}_{j}|}$ $\sum_{(x,y)\in D^{s}_{j}}{}\ell(\overline{\Phi^{t}}(x),y)$ Server’s loss corresponding to each partition}
@@
17:     Variance across server meta-losses: $\sigma_{D^{s}}^{2}(\overline{\Phi^{t}})=\frac{1}{|m|}\sum_{j\in m}(\mathcal{L}_{D^{s}_{j}}(\overline{\Phi^{t}})-\frac{1}{|m|}\mathcal{L}_{D^{s}}(\overline{\Phi^{t}}))^{2}$
18:     Customized server meta-loss: $\mathcal{L}_{meta}(\overline{\Phi^{t}})=\gamma\mathcal{L}_{D^{s}}(\overline{\Phi^{t}})+(1-\gamma)\sigma_{D^{s}}^{2}(\overline{\Phi^{t}})$
19:     $\Phi^{t}$leftarrow $\overline{\Phi^{t}}$- $\zeta\nabla\mathcal{L}_{meta}(\overline{\Phi^{t}})$ Model meta-updates using customized loss}
20:  end for
@@ C.1 Dataset Details
In the IID setup, data samples from each class are distributed equally to $|M|=100$ clients for pre-training and $|G|=10$ clients for downstream FL task. Taking the CIFAR-100 dataset in the IID pre-training phase as an example, there are $|M|=100$ clients, each holding 456 images. We will randomly select $|m|$ clients to participate in pre-training in each round. In the IID downstream phase of CIFAR-100, there are $|G|=10$ clients, each holding 300 images. In the non-IID setup, samples within each class are partitioned among $|M|$ and $|G|$ clients using a Dirichlet($\alpha$) distribution for pre-training and downstream task, respectively, with $\alpha=0.5$ selected as is in the literature (Morafah et al. 2022; Li, He, and Song 2021).
In addition to the label distributional shifts discussed in our manuscript, we also explore our method’s capability to handle domain shifts. To assess this, we conducted an additional experiment using the PACS dataset (Li et al. 2017). As PACS dataset consists of four domains (Art, Cartoon, Photo, Sketch), we separate them into three domains for pre-training and one domain for downstream FL, following the one-domain-leave-out setup (Li et al. 2022; Zhou et al. 2021).
@@ C.2 Hyperparameters and Compute Settings
We use ResNet-18 as the model structure for image classification, following the setting of (Nguyen et al. 2023; Chen et al. 2023). For our method, the SGD optimizer with a learning rate of $\eta=10^{-3}$ and $\zeta=10^{-3}$ is adopted for both local and meta updates. Both local and meta learning rates are searched within the range of [1e-2, 5e-3, 1e-3, 5e-4]. We searched for learning rates within the range of [1e-2, 5e-3, 1e-3, 5e-4] for local training of all FL pre-training baselines and selected 1e-3 as the optimal learning rate for them. In scenario II, each FL baseline will continue to conduct a few SGD iterations using the server’s data after constructing their global model. We searched for learning rates in the range of [1e-2, 1e-3] for this additional training and selected 1e-3 as the optimal learning rate for the server. Regarding hyperparameters in the q-FFL baseline, we conducted experiments with q-values of 1, 3, and 5 and reported the corresponding best statistics. We select a learning rate $\eta$ from the range [1e-2, 5e-3, 1e-3, 5e-4] for local updates in our CoPreFL and determined that 1e-3 provides the best results. Additionally, for meta-updates in both scenarios, we search for the learning rate $\zeta$ within the range [1e-2, 1e-3] and find that 1e-3 is the optimal value. In the case of the centralized baseline mentioned in Section 4.2, we searched for the optimal learning rate within the range [1e-2, 5e-3, 1e-3, 5e-4, 1e-4], ultimately selecting 1e-3. We utilized the SGD optimizer for all updates across all methods, and the batch size is set to be 32 for all experiments.
@@ D.1 Downstream FL Results with Scenario I Pre-training
Comparison of IID and non-IID downstream tasks. In some cases, when comparing IID and non-IID downstream tasks, we see that data heterogeneity has only a slight impact on the proposed approach. To be more specific, the performance of non-IID downstream FL experiences only a slight performance drop compared to IID downstream FL. In an IID setting, the model must classify all classes present in the system, as each client will have most classes in its local dataset. However, in a non-IID downstream scenario, the model may only need to classify a few classes since each client typically has a limited subset of classes. This simplification of local tasks in the non-IID setting reduces the complexity of evaluation for each client, minimizing the impact of data heterogeneity on the overall federated learning process. This phenomenon aligns with findings from other FL studies, where global model performance is assessed using local test sets from individual clients (Li et al. 2020; Diao, Li, and He 2023).
@@ D.2 Downstream FL Results with Scenario II Pre-training
Additional results with varying degrees of IIDness and numbers of participants. This section provides supplementary results for scenario II, where the server holds a small portion of the dataset during the pre-training phase. We also consider varying numbers of participants $|m|$, specifically 15, 20, 25, and 30 out of 100 clients, during the pre-training phase for these models. Tables 25, 26, 27, and 28 display the average performance across 10 IID FL downstream tasks and Tables 29, 30, 31, and 32 show the average performance across 10 non-IID FL downstream tasks. In both cases, the downstream FL were initialized by pre-trained models trained on 15, 20, 25, and 30 participants out of 100 clients, respectively, on the CIFAR-100 dataset. For the Tiny-ImageNet dataset, Tables 33, 34, 35, and 36 show the average performance across 10 IID FL downstream tasks and Tables 37, 38, 39, and 40 display the average performance across 10 non-IID FL downstream tasks. In both cases, the downstream FL were also initialized by pre-trained models trained on 15, 20, 25, and 30 participants out of 100 clients, respectively. It is important to note that in this scenario, FedAvg, FedMeta, and q-FFL undergo further training using server data through the SGD optimizer after each method completes its local iterations and obtains its respective global model in each round (Yang, Chen, and Shen 2023; Bian et al. 2023). Similarly, CoPreFL-SGD is trained using server data with the SGD optimizer on $\Phi^{t}$ in line 17 of Algorithm 1 in each round. This process involves conducting meta-updates and balancing performance and variance using clients’ data first, followed by updating the aggregate model again using the server’s dataset. Finally, CoPreFL follows Algorithm 2, utilizing server data for meta-updates. By incorporating meta-updates using server data to align with our objectives in (3), our pre-training method consistently outperforms other baselines, leading to improved average accuracy and reduced variance. Comparing CoPreFL with CoPreFL-SGD strongly suggests that, rather than conducting a few SGD iterations using server data, which may dilute our objectives, we recommend building pre-training objectives upon server data using meta-updates.
@@
Alternative implementation for baselines in scenario II. In addition to the hybrid training approach introduced in (Yang, Chen, and Shen 2023; Bian et al. 2023), which utilizes clients’ data ($\bigcup_{i\in M}D^{p}_{i}$) to train local models and then refines the aggregated global model using the server’s data ($D^{S}$), we explore an alternative implementation for other FL baselines in scenario II. In this case, we distribute the entire training dataset ($\bigcup_{i\in M}D^{p}_{i}+D^{S}$) to $|M|=100$ clients and select $|m|=20$ participants in each round for federated learning without a further refining step since there is no server’s data in this case. Therefore, each client holds more samples compared to their previous scheme. Table 41 shows the average performance of downstream non-IID FedAvg using non-IID federated methods as initialization. It is important to note that we maintain a fixed data splitting setup for our method, meaning we use $\bigcup_{i\in M}D^{p}_{i}$ for local training and meta-update the temporary model using $D^{S}$. These comparisons also show the superiority of our method.
@@ D.4 Details and Additional Results for FEMNIST Dataset
We also consider the FEMNIST dataset, widely used in FL research, following the data partition provided in (Park et al. 2021). We divide the 62 classes into 52 alphabet classes for the pre-training phase, reserving the remaining 10 digit classes for downstream FL tasks. Instead of using a ResNet-18 model, we employ a model consisting of two 3$\times$3 convolutional layers followed by two linear layers. We fixed the total number of clients as $|M|=100$ for pre-training and $|G|=10$ for downstream FL tasks. During the pre-training phase, we set the number of participants $|m|=20$ and the federated round $T=50$ for each federated pre-trained method. We use the SGD optimizer with a learning rate of $10^{-3}$ and batch size 32 for baselines and our method.
@@ D.5 Implementation Details and Additional Results for Different Initialization Methods
For FL baselines, we additionally consider SCAFFOLD (Karimireddy et al. 2019), which addresses partial client sampling, FedDyn (Acar et al. 2021), designed to tackle non-IID issues, and PerFedAvg (Fallah, Mokhtari, and Ozdaglar 2020), aiming to provide an adaptable personalized model, for a detailed comparison. We train all FL algorithms for 50 rounds with $|m|=20$ participants selected from $|M|=100$ clients under a non-IID setting (Dirichlet $\alpha=0.5$) , and the final global model is used as initialization for downstream FedAvg. In the case of non-IID related FL, FedDyn, we set the parameter $\alpha$ to 0.01. For personalized FL, PerFedAvg, we employ a two-step gradient descent for local client training introduced in their paper. We use the SGD optimizer with a learning rate of $10^{-3}$, a batch size of 32, and 50 federated rounds for these FL-based baselines.
@@ D.6 Implementation Details and Additional Results for Different Downstream FL Tasks
In addition to the general downstream FL tasks built by FedAvg, we consider FedProx (Sahu et al. 2018) and q-FFL (Li et al. 2020), more advanced FL algorithms that addresses heterogeneity and performance balance compared to FedAvg, to examine the robustness and generalizability of our pre-trained method. The experiments are conducted using the CIFAR-100 dataset under non-IID pre-training scenario I. Two additional FL algorithms, non-IID FedProx and non-IID q-FFL, are considered for downstream phase. We randomly sample 5 classes from the 20 available in our CIFAR-100 downstream dataset for each downstream task. The sampled data is then distributed to 10 clients, and the training consists of 50 rounds with 5 local iterations per round, utilizing an SGD optimizer with a learning rate of $10^{-3}$. We set the parameters $\mu=1$ for the proximal term coefficient in FedProx and $q=2$ for the loss-reweighting coefficient in q-FFL, following the optimal values reported by the authors for the CIFAR-100 dataset.
@@ D.7 Implementation Details when A Public Large-Scale Dataset is Available
In addition to utilizing CIFAR-100 and Tiny-ImageNet datasets, where we partition the datasets for pre-training and downstream tasks, we also explore a scenario where public large datasets are available for pre-training phase. We conducted experiments using pre-trained models with the ImageNet dataset (Deng et al. 2009), a widely used large public dataset for image classification. We sampled 200 images for each of the 1,000 classes in ImageNet_1K as a pre-training dataset. We conducted pre-training using both the centralized method and our proposed CoPreFL with ImageNet_1K. Subsequently, we conducted 10 non-IID FedAvg tasks using the CIFAR-100 dataset and initialized the models with these pre-trained models. For the centralized model in pre-training phase, we trained the model with the SGD optimizer and a learning rate of 1e-3, training the model for 50 epochs. For our proposed method during pre-training, we distributed all the sampled data across $|M|=100$ clients based on non-IID distribution (Dirichlet $\alpha=0.5$), sampling $|m|=20$ clients in each round, and conducted CoPreFL for 50 rounds. Since the goal of this experiment is to demonstrate that even with a centrally-stored public large dataset, we can intentionally distribute the dataset and apply our method, we only consider conducting our method under scenario I. For the downstream FL phase, we apply 10 IID and non-IID FedAvg tasks to the 20-class and 40-class downstream datasets we used in CIFAR-100 and Tiny-ImageNet, respectively. Each task executes FedAvg for $R=10$ rounds using the SGD optimizer with a learning rate of $10^{-3}$. Note that all classes observed during downstream tasks are the seen classes that have already appeared during pre-training, given that CIFAR-100 and Tiny-ImageNet are the subsets of ImageNet.